Suppose a class is containing 10 women and 7 men.
(a) In how many different ways can a team of 6 be chosen from the class?
(b) What if the team must contain exactly 4 women?
(c) What if the team must contain at least 4 women?

Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)


total 17: 10W, 7M
(a) In how many different ways can a team of 6 be chosen from the class? 17C6
(b) What if the team must contain exactly 4 women? (10C4)(7C2)
(c) What if the team must contain at least 4 women? (10C4)(7C2)+(10C5)(7C1)+(10C6)(7C0)

Answer by stanbon(75887)

Suppose a class is containing 10 women and 7 men.
(a) In how many different ways can a team of 6 be chosen from the class?
Ans: 17C6 = 12376 ways
-----------------------------------
(b) What if the team must contain exactly 4 women?
# of ways to succeed:: 10C4*7C2 = 4410
P(4 w and 2 m) = 4410/12376
-----------------------------------
(c) What if the team must contain at least 4 women?
# of ways to succeed 4410 + 10C5*7C1 + 10C6 = 6384
P(at least 4 w) = 6384/12376
------------------
Cheers,
Stan H.
-----------------